### v1

# Lifelong Graph Summarization with Neural Networks: 2012, 2022, and a Time Warp

Abstract.
Summarizing web graphs is challenging due to the heterogeneity of the modeled information and its changes over time. We investigate the use of neural networks for lifelong graph summarization. Assuming we observe the web graph at a certain time, we train the networks to summarize graph vertices. We apply this trained network to summarize the vertices of the changed graph at the next point in time. Subsequently, we continue training and evaluating the network to perform lifelong graph summarization. We use the GNNs Graph-MLP and GraphSAINT, as well as an MLP baseline, to summarize the temporal graphs. We compare $1$-hop and $2$-hop summaries. We investigate the impact of reusing parameters from a previous snapshot by measuring the backward and forward transfer and the forgetting rate of the neural networks. Our extensive experiments on ten weekly snapshots of a web graph with over $100$M edges, sampled in 2012 and 2022, show that all networks predominantly use $1$-hop information to determine the summary, even when performing $2$-hop summarization. Due to the heterogeneity of web graphs, in some snapshots, the $2$-hop summary produces over ten times more vertex summaries than the $1$-hop summary. When using the network trained on the last snapshot from 2012 and applying it to the first snapshot of 2022, we observe a strong drop in accuracy. We attribute this drop over the ten-year time warp to the strongly increased heterogeneity of the web graph in 2022.
The source code and additional resources: https://github.com/jofranky/Lifelong-Graph-Summarization-with-Neural-Networks
temporal graphs, lifelong graph learning, neural networks, graph neural networks, graph summary, RDF graph
1. Introduction
The idea of graph summarization is to generate a small representation $S$ of an input graph $G$, which preserves structural information necessary for a given task (Čebirić et al., 2019). Calculating the graph summary itself can be computationally expensive, especially for larger graphs (Blume et al., 2021). However, this reduction in size allows tasks, e.g., web data summary search (Konrath et al., 2012), entity retrieval (Ciglan et al., 2012), estimating cardinalities in database joins (Neumann and Moerkotte, 2011), and data visualization (Goasdoué et al., 2020), to be computed much faster than on the original graph (Blume et al., 2021). In machine learning, graph summaries are used to learn the embeddings for the original graph (Zhou et al., 2023) or for more efficient graph distillation (Gupta et al., 2023).
The information preserved by a summary is defined by the summary model and considers the vertices’ $k$-hop neighborhoods. Thus, in a machine learning sense, graph summarization can also be expressed as a vertex classification task where each vertex in the graph belongs to a certain summary (Blasi et al., 2022), representing the class. This summarization operation has to be permutation invariant to the input, i. e., the vertices’ $k$-hop neighborhood (Blume et al., 2021; Konrath et al., 2012). Graph Neural Networks (GNNs) are by design permutation invariant to the input (Hamilton, 2020) and thus can predict the summary of each vertex (Blasi et al., 2022).
Blasi et al. (Blasi et al., 2022) applied multiple GNNs to graph summarization. They considered graph summaries based on equivalence classes (Goasdoué et al., 2020) which are lossless with respect to specific features, such as the edge label of the outgoing edges of a vertex. The authors used the Dynamic Linked Data Observatory (DyLDO) dataset (Käfer et al., 2012) to compute the graph summary from the 6 May, 2012 snapshot. We extend the analysis of Blasi et al. by considering temporal graphs. We reuse the neural networks trained on snapshots from 2012 to summarize the snapshot from 2022. We call this a time warp.
The problem with training a network only on the first snapshot is that the performance decreases when the graph changes over time (Zhou and Cao, 2021; Wang et al., 2020). This is especially true for graph summaries, where classes may appear or disappear between snapshots (Blasi et al., 2022). One simple solution is learning the summaries for each snapshot from scratch, resulting in a separate network per snapshot without any information about summaries of previous snapshots. However, this is not feasible due to the size of the snapshots in the millions of vertices. In this work, we address this problem by incrementally learning the network. We reuse the network trained on previous snapshots as a base for learning the summaries on the next snapshot. This training procedure is called lifelong graph learning (Chen and Liu, 2018).
We first calculate the summaries of the vertices of each snapshot using a crisp algorithm (Blume et al., 2021). These summaries are used as the gold standard for training the GNNs. We continually train the networks on one snapshot after the other, i. e., we train one task after another. After training on a snapshot, we evaluate the network on the current, all past, and all future snapshots. We measure forward transfer, backward transfer, and forgetting rates of the networks based on classification accuracy (Lopez-Paz and Ranzato, 2017; Chaudhry et al., 2018). Forgetting is the performance drop of a network on snapshots trained on in earlier tasks after training on subsequent tasks (McCloskey and Cohen, 1989). Our results show that a network performs best on a snapshot $i$ when it is trained on the sequence of tasks up to that snapshot $i$. This observation is more prominent for the $2$-hop than the $1$-hop summary. After the ten-year time warp, reusing the network parameters from 2012 does not have a positive impact but also does not harm the performance in 2022. In summary, our contributions are:
  • We extend graph summarization using GNNs from static graphs to temporal graphs. We experiment with two GNNs and a baseline multi-layer perceptron (MLP) on sequences of ten weekly snapshots from 2012 and 2022 of the DyLDO web graph.
  • We show that a neural network performs best on a snapshot $i$ when being trained in a sequence of tasks up to and including that snapshot. The changes in the graph reduce the vertex classification performance. This observation is more prominent in 2022.
  • We show that an MLP is sufficient for a good classification performance in $2$-hop summaries, although it uses only $1$-hop information.
  • Reusing parameters from a network trained on the last snapshot in 2012 has no benefit over a network trained from scratch on the first snapshot in 2022.
The remainder of the paper is organized as follows: Below, we summarize related work. Section 3 introduces the neural networks and methods used for graph summarization. Section 4 describes the experimental apparatus. Our results are presented in Section 5. Section 6 discusses the results, before we conclude.
2. Related Work
First, we introduce the Graph Neural Networks (GNN) used in this work. Second, we look into recent approaches for graph summarization. Finally, we discuss literature on lifelong graph learning and pre-training.
2.1. Graph Neural Networks
A GNN uses the structural information of a graph contained in its edges and the features of the neighboring vertices to distinguish different vertices, e. g., to classify them. Many GNNs use a message-passing approach, e. g., the Graph Convolutional Network (GCN) (Kipf and Welling, 2017), the Graph Attention Network (Velickovic et al., 2018), and Graph Isomorphism Network (Xu et al., 2019).
GraphSAINT is a scalable GNN architecture using message-passing (Zeng et al., 2020). It is a sampling-based method that enables any GNN to be efficiently applied to large graphs. It uses a vertex-based, edge-based, or random-walk-based sampler to generate smaller subgraphs that are used to train the GNN.
Recently, Hu et al. (2021) introduced Graph-MLP, a GNN that does not use message passing. Graph-MLP consists of an MLP that is trained with a standard cross-entropy loss and a neighbor contrastive (NContrast) loss on the graph edges.
2.2. Graph Summarization
FLUID (Blume et al., 2021) is a language and a generic algorithm for flexibly defining and efficiently computing graph summaries. It can express all existing lossless structural graph summaries w.r.t. the considered features such as the vertices’ edges, the vertex neighbors, and others. This computation of graph summaries is based on hash functions applied on a canonical order of the vertices features (Blume et al., 2021; Konrath et al., 2012). The application of neural networks as a hash function has mainly been considered in the context of security (Turčaník and Javurek, 2016; Lian et al., 2007) but recently also applied by Blasi et al. with different GNN architectures for summarization of statics graphs (Blasi et al., 2022). In their work, Blasi et al. used three GNN architectures based on message-passing (GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), and GraphSAINT (Zeng et al., 2020)), Graph-MLP (Hu et al., 2021) as a GNN without message-passing, and Bloom filter. A sampling method based on GraphSAINT is applied that considers the class distribution by sampling inverse to the number of occurrences of classes in the training set.
The work on static graphs of Blasi et al. is the foundation of our work. We extend it to consider temporal graphs in a lifelong learning setting, which we introduce below. We reuse the sampling strategy from Blasi et al. for training our networks and use GCN as the underlying GNN.
2.3. Lifelong Graph Learning
Lifelong learning (Chen and Liu, 2018), also known as continual learning (Lesort et al., 2020), incremental learning (Chaudhry et al., 2018), and never-ending learning (Mitchell et al., 2018) is a learning procedure that imitates the lifelong learning ability of humans. It adds the importance of transferring and refining knowledge to the learned network. In lifelong learning (Fei et al., 2016), a network at time $t$ has performed a sequence of $t$ learning tasks $\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{t}$ and has accumulated the knowledge $\mathcal{K}$ learned in these past tasks. At time $t+1$, it is faced with a new learning task $\mathcal{T}_{t+1}$. The network can use past knowledge to help performing the learning task $\mathcal{T}_{t+1}$.
A lifelong learning model should be able to exploit previous knowledge to learn new tasks better or faster. Successful knowledge transfer occurs when the new tasks are similar to the previous ones (Ruvolo and Eaton, 2013). One issue of lifelong learning is catastrophic forgetting (McCloskey and Cohen, 1989). Catastrophic forgetting is the tendency of neural networks to abruptly and drastically forget previously learned information upon learning new information.
Lifelong learning is also applicable to graphs. In lifelong graph learning, a challenge is that the graphs can grow or shrink over time. Adding and removing vertices in the graphs may also result in the inclusion of new class labels in classification tasks (Zhou and Cao, 2021; Wang et al., 2020). The main challenges of lifelong graph learning are discussed in recent surveys (Zhang et al., 2024; Febrinanto et al., 2022). These are the ability to detect new instances of a class or tasks in the graph, the ability to accommodate new knowledge in graphs based on additional data instances or classes, and the ability to perform graph representation learning efficiently and accommodate new knowledge by sampling the graph. We address the aspect of increasing the detection accuracy of already-seen classes and adding new classes to our networks.
An issue with vertex classification problems is that the class of a vertex depends on its edges and, thus, is dependent on its neighbors. Wang et al. (2022) convert vertex classification tasks into graph classification tasks, i. e., turn the problem from lifelong graph learning into a task on i.i.d. data.
Our approach of lifelong graph learning follows the approach of Galke et al. (2023). The authors continue training the GNN on a new snapshot of the graph once it is available. Among others, they compare reusing an existing network (warm restart) versus training a new network from scratch for each task (cold restart). The results show that the warm restart benefits the lifelong training performance.
3. Graphs and Summarization
We define the basic notation of graphs used as input for our summary models. We introduce classical graph summarization, which is used to compute the gold standard, and graph summarization with neural networks. We describe our sampling approach and our training method. Finally, we describe how graph summarization is used in the lifelong learning setting.
3.1. Graphs
[…]
3.2. Classical Graph Summarization
A graph summary maps a graph $G$ to a smaller representation $S$, which is a graph that preserves structural information of $G$ (Scherp et al., 2023). Each vertex of the original graph $G$ is a member of exactly one equivalence class (EQC) of the summary graph. An EQC represents a set of vertices that share common features, such as the same set of edge labels. We formulate graph summarization as a vertex classification task following Blasi et al. (2022). Our summary models depend on the $k$-hop neighborhood of a vertex. The $1$-hop summary model is the Attribute Collection ($M_{AC1}$) that Blasi et al. (Blasi et al., 2022) used. The $2$-hop summary model $M_{AC2}$ extends the $1$-hop model by also considering the outgoing edges of the neighbors of a vertex. The models only make use of outgoing edges and only differ in the size of the neighborhood that is considered for the summary. In $ACi$, we calculate the EQC of a vertex $x$ by calculating $h_{i}(x)$ that considers the $i$-hop neighborhood of the vertex. We set $h_{1}(x)=0$ and then, for $i\geq 1$,
$h_{i+1}(x)=\bigoplus_{(p,h_{i}(y))\in H_{i}(x)}\mathrm{hash}(p,h_{i}(y))\,,$
where $H_{i}(x)=\{(p,h_{i}(y))\mid(x,p,y)\in E\}$ and $\bigoplus$ is the bitwise XOR operator. Note that $H_{i}(x)$ is a set and thus contains no duplicates. This is important since the XOR operation otherwise cancels out two values with identical hashes. The function $\mathrm{hash}(p,h_{i}(y))$ first concatenates $p$ and $h_{i}(y)$ to a string, which is then hashed. We use the default hashing function for strings in Python, which uses SipHash (Python, 2023; Aumasson and Bernstein, 2012).
An illustration of summarizing a vertex is shown in Figure 1. Our model considers, at most, the $2$-hop neighborhood of a vertex. In the case of vertex $v_{1}$, we recursively consider edges in the $2$-hop neighborhood of vertex $v_{1}$. Hence, we consider each edge in Figure 1 except the edge $(v_{6},p_{2},v_{9})$.
Figure 1. $3$-hop neighborhood of a vertex $v_{1}$.
We generalize this classical graph summarization by considering graphs that change over time. For a number of hops $i\in\{1,2\}$, the graph summary $S_{t}$ of a graph $G_{t}=(V_{t},E_{t},R_{t})$ at time $t$ is a tuple $(V^{S}_{t},E^{S}_{t},R^{S}_{t})$. $V^{S}_{t}=C\cup D$, where $C$ is the set of EQCs computed from the vertices $V_{t}$, and $D=\{d_{p,h_{i}(y)}\mid(x,p,y)\in E_{t}\text{ for some }x\}$ is a set of vertices further specifying the EQCs. These are the primary and secondary vertices of Blume et al. (Blume et al., 2021). $R^{S}_{t}$ is the set of predicates of the original graph $G_{t}$ that were considered for the summarization. $E^{S}_{t}$ are the outgoing edges of the summary vertices $V^{S}_{t}$, given by
$E^{S}_{t}=\{(v,(p,h_{i}(y)),d_{p,h_{i}(y)})\mid(x,p,y)\in E_{t}\text{ and }x\in C\}\,,$
where $d_{p,h_{i}(y)}\in D$.
3.3. Neural Graph Summarization
When applying graph neural networks (GNNs) for graph summarization, we need to define the vertex features (Hamilton, 2020; Blasi et al., 2022). Instead of hashing, as above, the feature vector $h_{v}^{(0)}$ of a vertex $v$ is a multi-hot encoding of its outgoing edge labels, following Blasi et al. (Blasi et al., 2022). The motivation is that the multi-hot encoded feature vector can be easily extended with new elements once new predicates appear in the temporal graph. We note that multi-hot encoding is appropriate for graph summarization, as graph summary models only depend on the existence of edge or vertex labels, not their multiplicity (Blume et al., 2021).
For the $1$-hop model $M_{AC1}$, the feature vector already contains all the information needed to calculate the summary, which is the label set of the vertex’s edges. For the $2$-hop model $M_{AC2}$, we consider the feature vector of $v_{1}$ and, in addition, consider the feature vectors of $v_{1}$’s neighbors. From these feature vectors, we know the labels of the outgoing edges of $v_{1}$ and its neighbors. Note that, in classical summarization, we also know which edge labels connect to which neighboring feature vectors. Here, we know the set of edge labels and the set of neighboring feature vectors but not which edge is associated with which neighbor. However, the neural networks can still differentiate between the equivalence classes in $M_{AC2}$.
A solution to incorporate the edges is to represent an edge between a vertex and its neighbor as a vertex between them, labeled with the edge label multi-hot encoded as a feature vector. In this case, we use two hidden layers in the GNN to cover the additional hop. Where this procedure is not applied, we use one hidden layer. We evaluate both approaches, i. e., test whether a ($k-1$)-layer message passing can approximate a $k$-hop summary.
For lifelong learning, we consider different snapshots of a graph $G$ associated with a timestamp $t$. At time $t$, we train a neural network to predict the summary $S_{t}$ of the graph $G$ at time $t$, which we refer to as task $\mathcal{T}_{t}$. We compute the number of EQCs present in the summary of time $t$ for each time $t$, and the number of added and deleted EQCs compared to time $t-1$. From this, we can also compute how many EQCs from time $t-1$ are still present at time $t$. We also count how many different EQCs in total have appeared in the snapshots $S_{1},\dots,S_{t}$, which corresponds to the size of the neural network’s output layer at time $t$.
3.4. Graph Neural Networks
We use two GNNs in our study, following Blasi et al. (Blasi et al., 2022): GraphSAINT (Zeng et al., 2020), a classical message-passing model designed to scale with large graphs, and Graph-MLP (Hu et al., 2021), which is an alternative to the usual message-passing GNNs.
GraphSAINT (Zeng et al., 2020) uses a sampler (vertex-, edge-, or random walk-based) to create a sampled graph $G_{P}=(V_{P},E_{P},R_{P})$ from a graph $G$. A GCN model is dimensioned to fit the sampled graph. The embedding $h^{k}_{v}$ of each vertex $v$ in layer $k$ is calculated as
$h^{k}_{v}=\sigma\left(\sum_{u\in V_{P}}\tilde{A}_{P}[v,u](W^{k-1})^{T}h_{u}^{(k-1)}\right),$
where $\tilde{A}_{P}$ is the normalized and sampled adjacency matrix, $W$ is the weight matrix, and $\sigma$ is the activation function. We use GraphSAINT with jumping knowledge connections and either one or two hidden layers (Xu et al., 2018).. In contrast to the original networks, we follow Blasi et al. (2022) and do not use layer normalization because the summary of a vertex depends on the structural information of its neighborhood, i. e., each edge is equally important. Additionally, we ran our experiments measuring the influence of layer normalization (see Appendix C), which comes to the same conclusion.
Graph-MLP (Hu et al., 2021) consists of a two-layer MLP followed by a classifier layer. The first layer ${X^{1}}=\text{Dropout}({\sigma}(XW^{0}))$ uses a GELU-activation function $\sigma$ and dropout. In the second layer $Z=X^{1}W^{1}$, the NContrast loss $\mathrm{loss}_{\mathrm{NC}}$ is applied. The output of the last layer $Y=ZW^{2}$ is used for the vertex classification, where standard cross entropy loss $\mathrm{loss}_{\mathrm{CE}}$ is applied. The NContrast loss of the $i$-th vertex is formulated as:
$\ell_{i}=-\log\frac{\sum^{B}_{j=1}1_{[j\neq i]}\gamma_{ij}\exp(\mathrm{sim}(z_{i},z_{j})/\tau)}{\sum^{B}_{k=1}1_{[k\neq i]}\exp(\mathrm{sim}(z_{i},z_{k})/\tau)},$
where $\mathrm{sim}(z_{i},z_{k})$ denotes cosine similarity. $z_{i}$ is the vertex embedding of the vertex to be classified after the second layer $Z$, and $z_{k}$ is the embedding of a sample vertex after the second layer $Z$. $\tau$ denotes the temperature parameter and $\gamma_{ij}$ is the strength of the connection between vertices $i$ and $j$.
$\gamma_{ij}\begin{cases}\neq 0&\text{if $j$ is in the $n$-hop neighborhood of }i\\ =0&\text{if $j$ is not in the $n$-hop neighborhood of }i\,.\end{cases}$
The NContrast loss of the $i$-th vertex considers $r$-hop neighbors of the vertex as positive samples, while the other vertices are negative samples. In our experiments, we only consider the $1$-hop neighbors of a vertex as positive samples. The total NContrast loss is then $\mathrm{loss}_{\mathrm{NC}}=\frac{1}{B}\sum^{B}_{i=1}\ell_{i}\,$. The combined loss is calculated as $\mathrm{loss}_{\mathrm{final}}=\mathrm{loss}_{\mathrm{CE}}+\alpha\cdot\mathrm{loss}_{\mathrm{NC}}\,$. The weighting coefficient $\alpha$ is used to balance the two losses. For $\alpha=0$, it is an MLP using only cross entropy as loss function.
3.5. Sampling Procedure
Due to the large size of our graphs, we use the sampler of Blasi et al. (2022) for training. The sampler is based on the GraphSAINT vertex sampler, with each vertex weighted according to the class distribution. This accounts for the imbalanced class distributions of the snapshots. The sampled subgraph of a vertex consists of either its $1$-hop neighborhood for $M_{AC1}$ or its $2$-hop neighborhood for $M_{AC2}$. The sampler is also applied to Graph-MLP to sample the $1$-hop neighborhoods. These are treated as positive samples. Vertices that are not connected via an edge are treated as negative samples.
One challenge of the sampling process when using the $2$-hop model is that the number of vertices can vary greatly between batches. We have to add all neighboring vertices of a sampled vertex since they determine the vertex’s class. However, the size of the neighborhood depends on the degrees of the sampled vertices. To avoid running out of memory, we restrict the sampler to a specific number of vertices in each batch as in Blasi et al. (2022).
4. Experimental Apparatus
In this section, we introduce our dataset and pre-processing, followed by a description of the experimental procedure. Afterward, we present how we optimize our hyperparameters. Finally, we present performance and forgetting measures for the evaluation of lifelong graph learning.
4.1. DyLDO Dataset in 2012 and 2022
The DyLDO dataset [^2] http://km.aifb.kit.edu/projects/dyldo/data/ was introduced by Käfer et al. (2012). It consists of weekly snapshots of the web. Starting from a seed list of about $90.000$ Uniform Resource Identifiers (URIs), it samples data from the web to create a snapshot. A snapshot consists line-based, plain text format for encoding Resource Description Framework (RDF) graphs (Gavin Carothers, 2014). The first snapshot was created on 6 May, 2012. The data contains some faulty and duplicate lines, which we removed following Blasi et al. (2022).
In our experiments, we use the following snapshots: The snapshots from 6 May, 2012 to 8 July, 2012 and the snapshots from 25 September, 2022 to 27 November, 2022. We have 20 snapshots in total, ten from each year. We use the ten snapshots from 2012 since that is when the sampling process started. The different numbers of EQCs in Figures 2(a) and 2(d) show that the dataset has much changed over the ten years. Thus, we use other snapshots from 2022, as well.
For each snapshot, we create the labels of the vertices by calculating their EQCs according to the summary model. We remove each vertex in a snapshot with a degree higher than $100$ for $M_{AC2}$ to be able to train on enough samples in a reasonable time for $M_{AC2}$. We also exclude them from the validation and test set. This has only a small effect on our dataset. For example, the number of EQCs removed for the summary $M_{AC2}$ of the first snapshot of DyLDO in 2012 is less than $5\%$, i. e., down from $288,418$ to $275,993$ EQCs.
4.2. Dataset Analysis
Figure 2 shows that the EQCs in each year change from snapshot to snapshot. We observe that the EQCs of the $2$-hop summary change more than the EQCs of the $1$-hop. Both changes underlie the same snapshot change, meaning that the $2$-hop summary changes more than the $1$-hop summary in the same sequence of snapshots. In general, there are more different EQCs in the $2$-hop summaries computed from the snapshots of 2022 than in 2012. The numbers of EQCs in the $1$-hop summaries are similar between the two years.
(a) May–July 2012
(b) $1$-hop, May–July 2012
(c) $2$-hop, May–July 2012
(d) September–November 2022
(e) $1$-hop, Sept.–Nov. 2022
(f) $2$-hop, Sept.–Nov. 2022
Figure 2. Detailed analyses of the DyLDO snapshots from 2012 and 2022 used in our experiments. Figures (a) and (d) show the absolute numbers of EQCs regarding the snapshots for a $1$-hop versus $2$-hop model in 2012 and 2022, respectively. The EQCs resemble the number of classes the neural networks are trained on. Figures (b), (c), (e), and (f) show the changes of the EQCs (addition, deletion, and recurring) in comparison to the previous snapshot.
4.3. Procedure
GNNs and baselines
We use a standard MLP with one hidden layer and ReLU-activation functions with dropout as a baseline. For $M_{AC1}$, we use an MLP because all information is contained in the feature vector. We have two cases for $M_{AC2}$, one with encoded edge information of the neighbors of a vertex and one without. In the first case, we only use GraphSAINT because the other neural networks do not use neighborhood information during inference. In the second case, we use all neural networks, i. e., GraphSAINT, MLP, and Graph-MLP. We train all networks with $100$ batches, each containing up to $1,000$ vertices (as described in Section 3.5), based on pre-experiments.
Experiments Summarizing DyLDO Snapshots from 2012 and 2022
We perform our experiments twice, once for our snapshots of $2012$ and once for our snapshots of $2022$. We incrementally train our network on each snapshot in chronological order. The vertices of each snapshot are divided into $2\%$ validation data and $5\%$ testing data. The remaining $93\%$ of the data is used for sampling the training data. Because each snapshot is very large, we can only train on parts of it.
For the first snapshot of each year, the neural networks are initialized with an input layer and an output layer with the size of the number of predicates and the number of classes, respectively. Then, the networks are trained on the snapshot for $100$ iterations with maximum batch size $1,000$. Each iteration has its own sampled batch. After training, we increase the input and output layer, if the next snapshot has more classes or predicates. The extended neural network is trained on the next snapshot for $100$ iterations. We repeat this process until we reach the last snapshot.
After we trained our networks, we tested each of the networks on all snapshots. We do this to measure the backward and forward transfer of the different networks (see Section 4.5).
Time Warp Experiment
In this experiment, we take the neural network trained on the 2012 snapshots and test it on the first snapshot in 2022. We contrast this reuse of a “10-year-old” network with a variant where the network is trained on the first snapshot of 2022. Finally, we compare these two “time warped” networks with a newly initialized network that is only trained on the first snapshot in 2022. This experiment shows whether reusing old networks of 2012 is helpful despite the time gap of ten years, compared to starting training from scratch in 2022.
4.4. Hyperparameter Optimization
For the hyperparameter search, we apply grid search on the parameter ranges from Blasi et al. (2022). We tune the learning rate on the values $\{0.1,0.01,0.001\}$, dropout in $\{0.0,0.2,0.5\}$, and hidden layer size $\{32,64\}$ for the GNNs. For Graph-MLP, we optimize the weighting coefficient $\alpha\in\{1.0,10.0,100.0\}$, the temperature $\tau\in\{0.5,1.0,2.0\}$, learning rate in $\{0.1,0.01\}$, and the hidden layer size $\{64,256\}$. For MLP, we use a hidden layer size of $1,024$ following prior works (Galke and Scherp, 2021; Blasi et al., 2022). We use Adam as an optimizer.
The best configurations depend on the summary model, namely $M_{AC1}$ and $M_{AC2}$. For the summary model $M_{AC1}$, we use an MLP with a hidden layer of size $1,024$, a dropout of $0.5$, and a learning rate of $0.01$. For $M_{AC2}$ with edge information, we use GraphSAINT with two hidden layers of size $32$, learning rate $0.1$, and no dropout. For $M_{AC2}$ with no edge information, we use an MLP with one hidden layer of size $1,024$, dropout $0.5$, and learning rate $0.01$. Furthermore, we use Graph-MLP with layer size $64$, dropout $0.2$, learning rate $0.01$, $\alpha=1$, and $\tau=2$. Finally, for GraphSAINT we use one hidden layer with hidden dimension $64$, learning rate $0.1$, and no dropout. Details can be found in Appendix A.
4.5. Measures
We assess vertex classification performance using accuracy. For lifelong learning, we base our analysis on the measures from Lopez-Paz and Ranzato (2017) and Chaudhry et al. (2018).
We train a network for each snapshot and test it on all other snapshots. Thus, for each of the years 2012 and 2022, we compute a result matrix $R\in\mathbb{R}^{T\times T}$ with $T=10$. Each entry $R_{i,j}$ is the test accuracy of the network on task $\mathcal{T}_{j}$ after observing the last sample from task $\mathcal{T}_{i}$. That means the network is trained on task $\mathcal{T}_{1}$ to $\mathcal{T}_{i}$ and tested on $\mathcal{T}_{j}$. We use these matrices to calculate the following measures for each year, to measure the backward and forward transfer of the networks. The average accuracy (ACC) a network achieves on all tasks $i$ after being trained on the last task $T$ is given by
$ACC=\frac{1}{T}\sum^{T}_{i=1}R_{T,i}\,.$
The backward transfer (BWT), which indicates how much knowledge of previous tasks the network keeps after being trained on the last task $T$ is given by
$BWT=\frac{1}{T-1}\sum^{T-1}_{i=1}R_{T,i}-R_{i,i}\,.$
The forward transfer (FWT), which shows how much knowledge of a task is reused for the next one, is defined as [^3] Different from Lopez-Paz and Ranzato (2017) who use the test accuracy $\bar{b}_{i}$ for the task $i$ at random initialization as a comparison, we use the performance $R_{i,i}$.
$FWT=\frac{1}{T-1}\sum_{i=2}^{T}R_{i-1,i}-R_{i,i}\,.$
We use the matrices to calculate further measures. The diagonal $R_{i,i}$ resembles the test accuracy for task $i$ after it is learned. Furthermore, $R_{i,1}$ is the test accuracy of the test data of the first task after $i$ tasks have been learned, i. e., the network is trained on the first $i$ snapshots and then tested in the first snapshot. Let $\alpha_{\mathrm{ideal}}=\max_{i\in\{1,\dots,T\}}R_{i,i}$ be the maximal test accuracy of the network trained up to task $i$ and then evaluated on task $i$. We define $\alpha_{i,\mathrm{all}}=\frac{1}{T}\sum^{T}_{t=1}R_{i,t}$ as the average accuracy of a network trained up to task $i$. With this, we can define the following measures (Kemker et al., 2018):
$\Omega_{\mathrm{base}}$ $=\frac{1}{T-1}\sum^{T}_{i=2}\frac{R_{i,1}}{\alpha_{\mathrm{ideal}}}$
$\Omega_{\mathrm{new}}$ $=\frac{1}{T-1}\sum^{T}_{i=2}R_{i,i}$
$\Omega_{\mathrm{all}}$ $=\frac{1}{T-1}\sum^{T}_{i=2}\frac{\alpha_{i,\mathrm{all}}}{\alpha_{\mathrm{ideal}}}\,.$
The first snapshot is the base test set, and the best accuracy $\alpha_{\mathrm{ideal}}$ in the matrix is the optimum. With $\Omega_{\mathrm{base}}$, we measure how a network performs on the base test set after training on other sets compared to the optimum. $\Omega_{\mathrm{new}}$ shows how a network performs on a snapshot after being trained on it. $\Omega_{\mathrm{all}}$ measures how a network performs on all previous snapshots compared to the optimum.
Finally, we use the accuracy matrix for computing the forgetting measure (Chaudhry et al., 2018):
$F_{k}=\frac{1}{k-1}\sum^{k-1}_{j=1}f_{j}^{k}\,,$
where $f_{j}^{k}$ is the forgetting on task $j$ after the network is trained up to task $k$ and is computed as:
$f_{j}^{k}=\max_{\ell\in\{1,\dots,k-1\}}R_{\ell,j}-R_{k,j}\,.$
5. Results
We describe the results for the vertex classification and analyze the lifelong learning measures. Afterward, we look into the forgetting of the networks. Finally, we investigate the ten-year time warp.
Vertex Classification
(a) MLP ($1$-hop)
(b) MLP ($2$-hop)
(c) Graph-MLP ($2$-hop)
(d) GraphSAINT ($2$-hop)
(e) GraphSAINT ($2$-hop and edges)
Figure 3. Accuracies for snapshots trained from May to July 2012.
(a) MLP ($1$-hop)
(b) MLP ($2$-hop)
(c) Graph-MLP ($2$-hop)
(d) GraphSAINT ($2$-hop)
(e) GraphSAINT ($2$-hop and edges)
Figure 4. Accuracies for snapshots trained from September to December 2022.
The classification accuracies for the snapshots in 2012 and 2022 are shown in Figures 3 and 4, respectively. In general, the networks perform well when they are tested on task $i$ after they were trained up to snapshot $i$.
Regarding 2012, the performance of the MLP for $M_{AC1}$ is much higher than for $M_{AC2}$. The difference between these two MLPs is between $0.25$ and $0.41$. Furthermore, the best performance of a network is observed when tested on the first snapshot independent of the training snapshots. For 2022, the results are similar to the experiments in 2012 but with overall lower accuracies. The MLP accuracies for $M_{AC1}$ are again much higher than those for $M_{AC2}$, but the difference is between $0.33$ and $0.53$. The best result for each snapshot is achieved by the network that has been trained on it. In contrast to 2012, the results on the first snapshot are not better than in the subsequent snapshots.
Table 1. Lifelong learning measures applied to the different network and summaries of each year.
Year Network Type ACC $\uparrow$ BWT $\uparrow$ FWT $\uparrow$ $\Omega_{\mathrm{base}}$uparrow $\Omega_{\mathrm{new}}$uparrow $\Omega_{\mathrm{all}}$uparrow
2012 MLP $1$-hop 0.739 -0.101 -0.062 0.898 0.829 0.817
2012 MLP $2$-hop 0.392 -0.087 -0.045 0.827 0.462 0.673
2012 Graph-MLP $2$-hop 0.420 -0.099 -0.049 0.831 0.499 0.675
2012 GraphSAINT $2$-hop 0.376 -0.087 -0.043 0.824 0.445 0.672
2012 GraphSAINT $2$-hop and edges 0.348 -0.098 -0.042 0.850 0.429 0.690
2022 MLP $1$-hop 0.701 -0.263 -0.275 0.715 0.938 0.654
2022 MLP $2$-hop 0.296 -0.289 -0.312 0.430 0.555 0.425
2022 Graph-MLP $2$-hop 0.290 -0.320 -0.311 0.437 0.573 0.426
2022 GraphSAINT $2$-hop 0.298 -0.276 -0.311 0.426 0.543 0.421
2022 GraphSAINT $2$-hop and edges 0.269 -0.273 -0.282 0.431 0.514 0.426
The lifelong learning measures applied to our measured accuracies are shown in Table 1. It shows that in 2012 and 2022, no network independent of the used summary model or information achieved a positive forward or backward transfer in our setting. For the other measures, the MLP applied on $M_{AC1}$ achieves the best results, and the results for $M_{AC2}$ are similar, independent of the network and the year.
Network Forgetting
Figure 5. The progress of the forgetting in 2012.
Figure 6. The progress of the forgetting in 2022.
The forgetting of the networks in 2012 is shown in Figure 5. The MLP ($1$-hop) has the highest forgetting, with over $0.13$, after being trained on the 10 June snapshot. Overall, the networks forget more after being trained on this snapshot. After the networks are trained on the last snapshot in 2012, they have a similar forgetting value.
The forgetting of the networks in 2022, as shown in Figure 6, differs from the ones in 2012. The highest forgetting for the $2$-hop networks is right in the beginning, and the lowest forgetting value is achieved by the MLP ($1$-hop) after being trained on the 16 October snapshot. In general, the forgetting values are higher in 2022 than in 2012. Nevertheless, after the networks are trained on the last snapshot, they have similar forgetting values again.
Ten-year Time Warp
Table 2. Accuracy after training on the task from September 25, 2022. Trained from scratch (no time warp, no TW) or initialized with weights from July 8, 2012 (/w TW).
Neural network Summary model /w TW no TW
MLP $1$-hop 0.933 0.930
MLP $2$-hop 0.581 0.572
Graph-MLP $2$-hop 0.605 0.612
GraphSAINT $2$-hop 0.565 0.572
GraphSAINT $2$-hop and edges 0.526 0.515
We show the results of reusing a neural network from 2012 to classify vertices in 2022 in Table 2. We call this setting a ten-year time warp (TW). Retraining the last network from 2012 on the first snapshot in 2022 gives the same results as training a network from scratch in 2022. Table 2 shows that the MLP with $M_{AC1}$ performs a bit better in the time warp. For $M_{AC2}$, the time warp makes the performance of the neural network better in two of four cases, and worse in the other cases. Reusing the last network from 2012 without retraining on the first 2022 snapshot resulted in very low accuracy – below $0.03$ in all cases (not shown in the table).
6. Discussion
Vertex Classification
Our experiments show that the problem of graph summarization by vertex classification is hard when applied to a temporal graph with changing and unseen classes. The results from the snapshots in $2012$ and $2022$ (Figures 3 and 4) show that the classification of EQCs for $M_{AC1}$ is easier than for $M_{AC2}$. This is caused by the increased number of EQCs when more hops are considered in the summary model. Figures 2(a) and 2(d) shows that there are more unique EQCs for $M_{AC2}$ than for $M_{AC1}$. The EQCs in $M_{AC2}$ also change more rapidly. For $2022$, the snapshots have more edges than the ones we use from $2012$. Overall, this reduces the performance. The more EQCs the network needs to learn and the more they change, the worse the network performs in general on all tasks.
Forward and Backward Transfer
We observe that no network has a positive forward or backward transfer between different snapshots. That means the networks perform better if they are trained on the task before being tested on it. The reason for the negative forward transfer is that new classes appear, and the reason for the negative backward transfer is that the neural network forgets classes. The results (Figures 3 and  4) confirm this statement, since a network always achieves its best accuracy on the snapshot it has been trained on. This observation is confirmed by $\Omega_{\mathrm{new}}$, the average of the accuracies on tasks after a network has been trained, which was higher than the other accuracies. One explanation is that the tasks the networks faced were very different, meaning that the EQCs of the snapshots change a lot, as can be seen in Figures 2(b), 2(c), 2(e), and 2(f).
The $\Omega_{\mathrm{base}}$ values in 2012 show that for this set of snapshots, the networks perform well on the first snapshot at each time. This is not the case for the networks for $M_{AC2}$ in 2022, where the values are below $0.5$. The networks do not perform as well as they could on the first snapshot. This shows that a network performs best on a task after being trained for it.
Forgetting and Reusing Networks
The $ACC$ and the $\Omega_{\mathrm{all}}$ value of a network indicate how well a v performs on previous tasks. Figures 5 and 6 show how much a network forgets over time. A higher $ACC$ and $\Omega_{\mathrm{all}}$ value do not necessarily mean a network forgets less. A good example is the MLP ($1$-hop) of 2012, which has a better $ACC$ and $\Omega_{\mathrm{all}}$ value than the $2$-hop models, but the progress of forgetting is higher.
Using $1$-hop versus $2$-hop Information in a $2$-hop Summary Model
We investigate whether the information provided from a $1$-hop neighborhood versus using the $2$-hop neighborhood is helpful. The measures show that the networks for $M_{AC2}$, independent of the used information, performed similarly (for both 2012 and 2022). Graph-MLP and GraphSAINT perform similarly to MLP for $M_{AC2}$, even though they have access to $2$-hop information for the classification task. While the $2$-hop information for GraphSAINT includes the path information between the vertices, the MLP can only use $1$-hop information. In summary, the results show that $2$-hop information does not improve the performance for a $2$-hop summary. While this seems to be counter-intuitive, an explanation is that for the $2$-hop summary, there are five to ten times more classes than in the $1$-hop summary since the number of paths grows exponentially with the number of hops (see Figures 2(a) and 2(d)). This results in many small EQCs, i. e., many classes with low support that are hard to train on (see Appendix B).
Warm versus Cold Restarts
Galke et al. (2023) differentiate between warm restart and cold restart in the context of lifelong learning. A warm restart is reusing network parameters from previous tasks, and retraining from scratch is called a cold restart. Our time warp experiment (Table 2) does not show a clear benefit of a warm restart compared to a cold restart over a ten-year range. In general, it shows that a warm restart does not hurt a network’s performance.
Limitations and Threat to Validity
Our networks may perform better if we trained them for more iterations. However, in Appendix A, we increased the number of training iterations by an order of magnitude and observed that it has only a small effect of less than $2$ points on the classification accuracy. If we continue training on the current task it may put the networks at more risk of forgetting the previous tasks. Furthermore, we use a hash function to compute the gold standard of the vertex labels, i. e., the EQCs. This could lead to a collision in the hash function. However, we chose the hash such that the probability is extremely low following Blasi et al. (Blasi et al., 2022).
7. Conclusion and Future Work
We show that graph summarization for temporal graphs using a neural network is more challenging than the equivalent task on static graphs. The performance of the vertex classification in a temporal graph depends on the number of classes and the changes between timesteps. A problem with the summarization is that new EQCs appear and known ones disappear over time. We observe forgetting in lifelong graph summarization by vertex classification. We argue that the performance of a network should not only be measured with the accuracy of the current task but also other measures should be taken into account, like forward and backward transfer. The lifelong learning measures show that the networks for $M_{AC2}$ perform similarly. Our experiments also show that just reusing the network parameters does not necessarily improve the lifelong learning measures.
In future work, one may use different lifelong learning approaches to mitigate forgetting and increase backward and forward transfer for lifelong graph summarization. It may be interesting to investigate different summaries and the impact of the singleton EQCs, i. e., where only a single vertex is summarized, which we did not address here. It may also be interesting to measure the similarity of the predicted EQCs to the actual EQCs computed from a graph summary algorithm. In our setting, retraining is always the best option, but future work can also look into the question of when to retrain, as done by Galke et al. (Galke et al., 2023).
Acknowledgements.
This research is co-funded by the CodeInspector project (No. 504226141) of the DFG, German Research Foundation. The authors acknowledge support by the state of Baden-Württemberg through bwHPC.
Appendix A Hyperparameter Optimization
The results of our hyperparameter optimization are shown in Table 3. Since the search is very expensive, we first perform a grid search with each network and configuration for $46$ iterations. Subsequently, we narrowed the search and chose the three best configurations of a network for each summary. We validate every neural network for $751$ more iterations, except MLP when using the graph summary model $M_{AC1}$. For MLP with $M_{AC1}$, we use only $121$ iterations because its batches contain more samples than for the graph neural networks. The results show that training for more iterations does not improve the performance.
Network Model LS LR D $\alpha$ $\tau$ Acc(1) Acc(2) Acc(3) Acc(4) Acc(5)
MLP $M_{AC1}$ 1,024 0.1 0.5 0.8481 0.8380 0.8357 0.8297 0.8440
MLP $M_{AC1}$ 1,024 0.01 0.5 0.7969 0.8398 0.8499 0.8496 0.8532
MLP $M_{AC1}$ 1,024 0.1 0.2 0.8378 0.8367 0.8293 0.8366 0.8367
MLP $M_{AC2}$ 1,024 0.01 0.5 0.5758 0.5775 0.5807 0.5832 0.5752
MLP $M_{AC2}$ 1,024 0.1 0.5 0.5479 0.5562 0.5524 0.5548 0.5458
MLP $M_{AC2}$ 1,024 0.1 0.2 0.5539 0.5633 0.5546 0.5543 0.5518
GraphSAINT $M_{AC2}$ 64 0.1 0.2 0.5654 0.5589 0.5714 0.5575 0.5621
GraphSAINT $M_{AC2}$ 32 0.1 0.2 0.5686 0.5639 0.5518 0.5559 0.5588
GraphSAINT $M_{AC2}$ 64 0.1 0 0.5736 0.5797 0.5720 0.5788 0.5686
GraphSAINT (Edges) $M_{AC2}$ 32 0.1 0 0.5558 0.5665 0.5550 0.5656 0.5540
GraphSAINT (Edges) $M_{AC2}$ 64 0.1 0 0.5409 0.5566 0.5481 0.5389 0.5352
GraphSAINT (Edges) $M_{AC2}$ 64 0.1 0.5 0.5350 0.5437 0.5223 0.5350 0.5342
Graph-MLP $M_{AC2}$ 64 0.01 0.2 1 2 0.6184 0.6250 0.6280 0.6180 0.6286
Graph-MLP $M_{AC2}$ 64 0.01 0 1 0.5 0.6222 0.6294 0.6253 0.6257 0.6269
Graph-MLP $M_{AC2}$ 64 0.01 0 10 0.5 0.6195 0.6228 0.6210 0.6246 0.6258
Table 3. Extensive search for parameters for the networks. Abbreviations: Layer Size (LS), Learning Rate (LR), and Dropout (D)). For MLP ($M_{AC1}$), the accuracy Acc($n$) is measured after the network has been trained for $1+n\cdot 24$ iterations and for the other networks after they have been trained for $1+n\cdot 150$ iterations. Best networks are in bold.
Appendix B Graph Summary Measures
To analyze the graphs and summaries more extensively, we exploit additional measures. We divide these measures into three categories. The first category is the unary measures. The second is binary measures, and the last is meta-measures.
B.1. Unary Measures
Unary measures only consider a summary $S_{t}$ at time $t$. The features of each vertex in the summary are the hash value that indicates its EQC, the degree, and the members of the EQC in the original graph. For each summary $S_{t}$, we have a function $ext_{t}()$ that returns for each vertex in the summary the number of vertices extensions in the original graph $G_{t}$. For the vertices, i.e., EQCs in the summary, $S_{t}$, we calculate the average size of EQCs as
(1) $\delta^{(1)}_{AvgSize}(S_{t})=\frac{\sum_{q\in V^{S}_{t}}ext_{t}(q)}{|V^{S}_{t}|},$
and the average number of edges/attributes per EQC as
(2) $\delta^{(1)}_{AvgE}(S_{t})=\frac{|E^{S}_{t}|}{|V^{S}_{t}|}.$
We also calculate distributions over the number of attributes of an EQC, over the number of members of an EQC, and over the number of usages of a predicate.
Figure 7 shows the average sizes of the EQCs as well as the average numbers of edges. Example distributions are depicted in Figures 8 and 9 for one snapshot in 2012 and 2022, respectively.
(a) Average number of attributes per EQC in 2012 ($1$-hop).
(b) Average number of members per EQC in 2012 ($1$-hop).
(c) Average number of attributes per EQC in 2012 ($2$-hop).
(d) Average number of members per EQC in 2012 ($1$-hop).
(e) Average number of attributes per EQC in 2022 ($1$-hop).
(f) Average number of members per EQC in 2022 ($1$-hop).
(g) Average number of attributes per EQC in 2022 ($2$-hop).
(h) Average number of members per EQC in 2022 ($2$-hop).
Figure 7. Average numbers for attributes (AvgE) and members (AvgSize) per EQC in the snapshots 2012 and 2022.
(a) Frequency of number of attributes per EQC (in descending order) for the $1$-hop attribute collection.
(b) Frequency of number of attributes per EQC (in descending order) for the $2$-hop attribute collection.
(c) Frequency of number of members of an EQCs (in descending order) for the $1$-hop attribute collection.
(d) Frequency of number of members of an EQCs (in descending order) for the $2$-hop attribute collection.
(e) Frequency how often a predicate is used in the original graph (in descending order).
Figure 8. Distributions measures on the summary of 2012-05-06
(a) Frequency of number of attributes per EQC (in descending order) for the $1$-hop attribute collection.
(b) Frequency of number of attributes per EQC (in descending order) for the $2$-hop attribute collection.
(c) Frequency of number of members of an EQCs (in descending order) for the $1$-hop attribute collection.
(d) Frequency of number of members of an EQCs (in descending order) for the $2$-hop attribute collection.
(e) Frequency how often a predicate is used in the original graph (in descending order).
Figure 9. Distributions measures on the summary of 2022-09-25
B.2. Binary Measures
Binary measures are used to compare a current summary $S_{t+1}$ to the previous summary $S_{t}$. It is used to measure the change that occurred between $S_{t+1}$ and $S_{t}$. The following measure is calculated with the Jaccard coefficient to calculate the change of vertices:
(3) $\delta^{(2)}_{JacV}(S_{t},S_{t+1})=1-\frac{|V^{S}_{t}\cap V^{S}_{t+1}|}{|V^{S}_{t}\cup V^{S}_{t+1}|}.$
We also calculate the Jensen–Shannon Divergence of $S_{t+1}$ and $S_{t}$. The Jensen–Shannon Divergence can be seen as the relative entropy between two probability distributions and is calculated as
(4) $\delta^{(2)}_{JS}(A,B)=D(A,B)+D(B,A)$
with
(5) $D(A,B)=\sum_{q\in V^{B}}P_{A}(q)\log_{2}\frac{P_{A}(q)}{P_{B}(q)}$
and $P_{A}(q)=\frac{ext_{t}(q)}{|V_{A}|},$ where $V_{A}$ are the vertices in A. If $ext_{t}(q)=0$, then $P_{A}(eq)=0$, and we consider the summand to be 0 as
$\lim_{x\rightarrow 0}x\log(x)=0.$
The changes in the binary measures are shown in Figure 10.
B.3. Meta-Measures
With meta-measures, we track how the summaries change over time, and in contrast to binary measures, it considers more than two summaries. During our experiments, we track whether new EQCs are appearing compared to the first snapshot and the previous one. We also track whether EQCs are reappearing or disappearing and how many EQCs have been seen until a certain snapshot. The meta-measures of our snapshots in 2012 and 2022 are depicted in Figure 11.
(a) Jaccard Coefficients in 2012.
(b) Jaccard Coefficients in 2022.
(c) Jensen-Shannon Divergences in 2012.
(d) Jensen-Shannon Divergences in 2022.
Figure 10. Jaccard Coefficients and Jensen-Shannon Divergences in the snapshots of 2012 and 2022.
(a) Changes of the EQCs compared to the first one in 2012 ($1$-hop).
(b) Changes of the EQCs compared to the previous one in 2012 ($1$-hop).
(c) Changes of the EQCs compared to the first one in 2012 ($2$-hop).
(d) Changes of the EQCs compared to the previous one in 2012 ($2$-hop).
(e) Changes of the EQCs compared to the first one in 2022 ($1$-hop).
(f) Changes of the EQCs compared to the previous one in 2022 ($1$-hop).
(g) Changes of the EQCs compared to the first one in 2022 ($2$-hop).
(h) Changes of the EQCs compared to the previous one in 2022 ($2$-hop).
(i) Number of EQCs in each snapshot of 2012.
(j) Number of all seen EQCs till a snapshot in 2012.
(k) Number of EQCs in each snapshot of 2022.
(l) Number of all seen EQCs till a snapshot in 2022.
Figure 11. Measuring the changes in the EQCs per snapshot in 2012 and 2022.
Appendix C Impact of Layer Normalization
We investigate the impact of normalizing a GCN, we normalize the GCN by using the degrees of the vertices. The adapted activation $h^{k}_{v}$ of each vertex $v$ in layer $k$ is calculated as follows:
(6) $h^{k}_{v}=\sigma\left(\sum_{u\in V_{s}}\frac{1}{\sqrt{\hat{d}_{v}\hat{d}_{u}}}\tilde{A}_{P}[v,u](W^{k-1})^{T}h_{u}^{(k-1)}\right)\,,$
where $\hat{d}_{v}$ is the degree of $v$ and $\hat{d}_{u}$ is the degree of $u$.
The results of the accuracy scores depicted in Figure 12 show that normalization does not make a big difference. This also is true of the forgetting scores, which are reported in Tables 4 and 5.
(a) GraphSAINT ($2$-hop)
(b) GraphSAINT ($2$-hop and normalized).
Figure 12. Accuracies comparison (Normalization vs. no normalization)
Year Network Type ACC $\uparrow$ BWT $\uparrow$ FWT $\uparrow$ $\Omega_{\mathrm{base}}$uparrow $\Omega_{\mathrm{new}}$uparrow $\Omega_{\mathrm{all}}$uparrow
2012 GraphSAINT $2$-hop 0.376 -0.087 -0.043 0.824 0.445 0.672
2012 GraphSAINT $2$-hop and normalized 0.372 -0.091 -0.042 0.825 0.445 0.675
Table 4. Lifelong Learning Measures (Normalization vs. no normalization).
Year Network Type $F_{2}$ $F_{3}$ $F_{4}$ $F_{5}$ $F_{6}$ $F_{7}$ $F_{8}$ $F_{9}$ $F_{10}$
2012 GraphSAINT $2$-hop 0.010 0.042 0.071 0.032 0.037 0.080 0.044 0.071 0.087
2012 GraphSAINT $2$-hop and normalized 0.019 0.047 0.064 0.017 0.047 0.081 0.049 0.065 0.093
Table 5. Forgetting values (Normalization vs. no normalization).